# Performance Estimation in Binary Classification Using Calibrated Confidence

### v1

@@ D.2 Simulated covariate shift
For both test sets, we ran 1,000 trials with a monitoring window size of 1,000 as well. In each trial, we drew a random sample from the test set, estimated the performance using the four metrics of our interest, and compared those against the actual values of the metrics over the whole test set. The results are shown in Table 4, where for both test sets we report class imbalance (the prevalence of the positive class) as "Bias", calibration error of each model as "ACE", the actual values for the four metrics of interest, and the mean absolute estimation errors for the CBPE method as "MAE". The error bounds for the estimation errors are given in the form of two standard deviations. All values are percentages except error bounds, which are given as percentage units.
@@
Number of Accuracy Precision Recall F1
Features Split Bias ACE True MAE True MAE True MAE True MAE
2 ID 32.7 0.71 87.5 -0.16$\pm$12 84.9 1.35$\pm$12 75.2 -2.44$\pm$12 79.8 -0.83$\pm$12
OoD 43.8 2.40 71.7 2.08$\pm$11 69.0 5.00$\pm$11 64.1 -3.65$\pm$11 66.5 0.07$\pm$11
10 ID 40.1 0.84 83.2 0.24$\pm$5.7 80.5 0.90$\pm$5.7 76.9 -0.66$\pm$5.7 78.7 0.07$\pm$5.7
OoD 48.0 1.03 76.1 0.12$\pm$3.9 75.7 0.95$\pm$3.9 73.9 -1.59$\pm$3.9 74.8 -0.38$\pm$3.9
100 ID 40.2 1.04 83.4 0.03$\pm$4.3 80.2 0.13$\pm$4.3 77.9 0.23$\pm$4.3 79.0 0.17$\pm$4.3
OoD 47.7 1.31 75.7 1.09$\pm$2.2 75.3 0.56$\pm$2.2 73.1 1.85$\pm$2.2 74.2 1.21$\pm$2.2
1000 ID 40.1 0.74 83.3 -0.30$\pm$4.9 80.2 -0.14$\pm$4.9 77.5 -0.75$\pm$4.9 78.9 -0.47$\pm$4.
OoD 47.8 1.77 74.6 1.74$\pm$2.5 73.9 1.78$\pm$2.5 72.4 1.39$\pm$2.5 73.2 1.58$\pm$2.5

### v2

@@ D.1 Simulated covariate shift
For both test sets, we ran 1,000 trials with a monitoring window size of 1,000 as well. In each trial, we drew a random sample from the test set, estimated the performance using the four metrics of our interest, and compared those against the actual values of the metrics over the whole test set. The results are shown in Table 4, where for both test sets we report class imbalance (the prevalence of the positive class) as ”Bias”, calibration error of each model as ”ACE”, the actual values for the four metrics of interest, and the mean absolute estimation errors for the CBPE method as ”MAE”. The error bounds for the estimation errors are given in the form of two standard deviations. All values are percentages except error bounds, which are given as percentage units.
@@
Fea- Accuracy Precision Recall F1
tures Split Bias ACE True MAE True MAE True MAE True MAE
[…]